OpenStudy (anonymous):

find the product (4x+3y)^2

OpenStudy (anonymous):

16x^2 +24xy +9y^2

OpenStudy (anonymous):

(4x)^2 + (3y)^2 + 2(4x)(3y) = 16x^2+9y^2+24xy since (a+b)^2 = a^2+b^2+2ab